Calculating the probability of picking out 3 pieces of the same kind of candy. The bag contains 6 pieces of chocolate, 4 pieces of fruity, 2 pieces of mint. Please help! Thanks.
Answer by stanbon(75887)

Calculating the probability of picking out 3 pieces of the same kind of candy. The bag contains 6 pieces of chocolate, 4 pieces of fruity, 2 pieces of mint.
=====
# of ways to succeed:
3 chocolate or 3 fruity = 6C3 + 4C3 = 20+4 = 24 ways
----
# of possible outcomes: 12C3 = 220 ways
-----
P(3 of same) = 24/220 = 12/110 = 6/55
=============
Cheers,
Stan H.
============